(9.19) A flywheel turns at 480 rpm. Compute the angular speed at any point on the wheel and
the tangential speed if it has a radius of 30 cm. Note: tangential speed can be considered as the
linear speed of the wheel.
1
Expert's answer
2021-07-13T13:32:23-0400

rpm=480

to convert rpm to rad/s

"rad/s=rpm\times2\frac{\pi}{60}"

"=480\times2\frac{\pi}{60}"

"=50.29rad\/s"


tangential speed

V"_{r} =r\omega"

radius is 30cm which is 0.3m

"50.29\times0.3m=15.1"

V"_{r} =15.1m\/s"